OpenStudy (anonymous):

Choose the equation of the line passing through the point (-2, 6) and parallel to y = 1/2 x - 8.

OpenStudy (ash2326):

the given line is y=1/2x-8 its slope is 1/2 so line parallel to it will have slope 1/2 y=mx+c m=1/2 c=? y=1/2x+c line passes through -2,6 so this will satisfy this line's equation 6=-1+c c=7 y=1/2x+7 solution